# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

### v3

@@ 10. Conclusion
All experiments were conducted under strict ethical guidelines to ensure data privacy and experimental safety. The evaluation relied exclusively on publicly available datasets, including ASVSpoof 2019 for benchmarking spoofing countermeasures (Todisco2019ASVspoof) and LibriSpeech for end-to-end system evaluation (panayotov2015librispeech). No personal data was collected from real-world victims. Additionally, assessments involving commercial APIs (e.g., Microsoft Azure (microsoft_speaker)) were performed using these standardized datasets in controlled environments, ensuring no disruption to live services or unauthorized access to legitimate user accounts.

### v4

@@ 10. Conclusion
All experiments were conducted under strict ethical guidelines to ensure data privacy and experimental safety. The evaluation relied exclusively on publicly available datasets, including ASVSpoof 2019 for benchmarking spoofing countermeasures (et al., 2019) and LibriSpeech for end-to-end system evaluation (Panayotov et al., 2015). No personal data was collected from real-world victims. Additionally, assessments involving commercial APIs (e.g., Microsoft Azure (Microsoft, )) were performed using these standardized datasets in controlled environments, ensuring no disruption to live services or unauthorized access to legitimate user accounts.